# Charactarisation of distal actions of automorphisms on the space of one-parameter subgroups of Lie groups

### v1

@@ 1. Introduction
For a Hausdorff topological space, a homeomorphism $T$ of $X$ is said to be distal (equivalently, $T$ acts distally on $X$) if for any pair of distinct elements $x,y\in X$, the closure of the double orbit $\{(T^{n}(x),T^{n}(y))\mid n\in{\mathbb{Z}}\}$ in $X\times X$ does not intersect the diagonal, i.e. for $x,y\in X$ with $x\neq y$, $\overline{\{(T^{n}(x),T^{n}(y))\mid n\in{\mathbb{Z}}\}}\cap\{(a,a)\mid a\in X\}=\emptyset$. David Hilbert introduced the notion of distality to study non-ergodic actions on compact spaces (see Ellis [16] and Moore [29]). Distal actions on compact spaces, Lie groups as well as locally compact groups have been studied by many mathematicians in different contexts (see Abels [1, 2], Choudhuri et al [10], Ellis [16], Furstenberg [17], Palit-Shah [30], Palit-Prajapati-Shah [31], Raja-Shah [33, 34], Shah [36], Shah-Yadav [38, 39, 40], and references cited therein).
@@
Let $G$ be a locally compact (Hausdorff) group with the identity $e$ and let ${\rm Sub}_{G}$ denote the set of all closed subgroups of $G$ endowed with the Chabauty topology. Note that ${\rm Sub}_{G}$ is compact and Hausdorff, and it is metrisable if $G$ is second countable [6]. Let ${\rm Aut}(G)$ denote the group of all (bicontinuous) automorphisms of $G$. There is a canonical group action of ${\rm Aut}(G)$ on ${\rm Sub}_{G}$; namely, $(T,H)\mapsto T(H)$, $T\in{\rm Aut}(G)$, $H\in{\rm Sub}_{G}$. Equivalently, we have a homomorphism from ${\rm Aut}(G)$ to ${\rm Homeo}({\rm Sub}_{G})$. As the image of ${\rm Aut}(G)$ under this action is a large subclass of homeomorphisms of ${\rm Sub}_{G}$, it is important to study the dynamics of the action of this special subclass in terms of distality. In [40], Shah and Yadav have studied the action of automorphisms of a connected Lie groups $G$ on ${\rm Sub}_{G}$ extensively and have shown in particular that for a large class of $G$ without the compact central subgroup of positive dimension, an automorphism $T$ of $G$ acts distally on ${\rm Sub}^{a}_{G}$, the space of closed abelian subgroups of $G$ if and only if $T$ generates a relatively compact subgroup of ${\rm Aut}(G)$. A similar characterisation also holds for $T$ belonging to the class (NC), i.e. those $T$ for which the closure of the $T$-orbit of any discrete cyclic group does not contain the trivial group $\{e\}$. Note that ${\rm Sub}^{a}_{G}$ is very large in many groups, e.g. it is the same as ${\rm Sub}_{G}$ if $G$ is abelian. It also contains all discrete cyclic subgroups. Shah, with Palit in [30], and with Palit and Prajapati in [31], has studied distal actions of automorphisms of a discrete group $G$ on ${\rm Sub}^{a}_{G}$, where $G$ is either polycyclic or a lattice in a connected Lie group.
Here, we would like to consider a closed subset of ${\rm Sub}^{a}_{G}$ containing smallest closed connected subgroups of $G$; namely, the closure in ${\rm Sub}_{G}$ of the set of all closed one-parameter subgroups of $G$; we denote this closed set by ${\rm Sub}^{p}_{G}$. Note that ${\rm Sub}^{p}_{G}$ is invariant under the action of ${\rm Aut}(G)$. We would like to study the dynamics of the actions of automorphisms on ${\rm Sub}^{p}_{G}$ and characterise them in terms of distality. In the case of ${\mathbb{R}}^{n}$, $n\geq 2$, ${\rm Sub}^{p}_{{\mathbb{R}}^{n}}$ is homeomorphic to ${\mathbb{RP}}^{n-1}\sqcup\{\{0\}\}$. It is indeed much smaller than ${\rm Sub}^{a}_{{\mathbb{R}}^{n}}={\rm Sub}_{{\mathbb{R}}^{n}}$. In general for a Lie group $G$, ${\rm Sub}^{p}_{G}$ does not contain any non-trivial discrete subgroup as it can not be a limit point of closed one-parameter subgroups (cf. Lemma 2.2). Theorem 4.1 of Shah and Yadav in [40] mainly deals with the action of automorphisms on discrete cyclic subgroups. Here, we are concentrating on the action of $T$ on ${\rm Sub}^{p}_{G}$, the smallest compact subset of ${\rm Sub}_{G}$ which contains all closed (connected) one-parameter subgroups of $G$. and get a characterisation for the distal action for a large class of connected non-abelian Lie groups $G$.
@@
Note that the condition on the center in Theorem 1.1 holds if $G$ is simply connected, or more generally its nilradical is simply connected. It also holds for $G/M$ for any connected Lie group $G$, where $M$ is the maximal central torus in $G$ (the maximal compact connected central subgroup in $G$). The result in case of non-abelian groups $G$ in (2) as above gives a similar characterisation as in [40], even though the space ${\rm Sub}^{p}_{G}$ is much smaller than ${\rm Sub}^{a}_{G}$. In case $G$ as above is abelian, then $G={\mathbb{R}}^{d}$, a vector group, and ${\rm Sub}^{p}_{G}$ is homeomorphic to ${\mathbb{RP}}^{d-1}\sqcup\{\{0\}\}$ as noted earlier, and the Theorem 1.1 (1) shows that a larger class of $T$ act distally on ${\rm Sub}^{p}_{G}$.
In [40], it is shown for a connected Lie group without any compact central subgroup of positive dimension that if $T$ acts distally on ${\rm Sub}^{a}_{G}$, then it acts distally on $G$ (more generally, see Theorem 3.6 and Corollary 3.7 of [40]). We generalise this for Lie groups which are not vector groups; as in the later case, Theorem 1.1 (1) shows that it is possible to have automorphisms which act distally on ${\rm Sub}^{p}_{G}$ but not on $G$.
@@
Theorem 1.3 illustrates in particular that the converse of Theorem 1.2 does not hold as the non-trivial unipotent automorphisms do not act distally on ${\rm Sub}^{p}_{G}$.
Let ${\rm Sub}^{c}_{G}$ denote the set of all closed discrete cyclic subgroups of $G$. It is invariant under the action of ${\rm Aut}(G)$ and so is its closure in ${\rm Sub}_{G}$. In [30] and [31], distal actions of automorphisms of ${\rm Sub}^{c}_{\Gamma}$ have been characterised, where $\Gamma$ is a certain type of discrete group, e.g. a discrete polycyclic group or a lattice in a connected Lie group. In a Lie group $G$, any closed one-parameter subgroup $H=\{x_{t}\}_{t\in{\mathbb{R}}}$ is the limit of $\{H_{n}\}_{n\in{\mathbb{N}}}$ as $n\to\infty$, where $H_{n}=\{x_{m/n}\mid m\in{\mathbb{Z}}\}$. Here, $H_{n}$ is a closed discrete cyclic group generated by $x_{1/n}$. If $H$ is compact, we may assume that $x_{1}=e$, the identity in $G$. Thus ${\rm Sub}^{p}_{G}\subset\overline{{\rm Sub}^{c}_{G}}$ (the closure of ${\rm Sub}^{c}_{G}$ in ${\rm Sub}_{G})$ and the latter is contained in ${\rm Sub}^{a}_{G}$. We should note that $\overline{{\rm Sub}^{c}_{G}}$ is much bigger than ${\rm Sub}^{p}_{G}$ as the latter contains all discrete cyclic subgroups, and these do not belong to ${\rm Sub}^{p}_{G}$ (see Lemma 2.2). It is easy to see that Theorems 1.1 (2), 1.2 and 1.3 remain valid if ${\rm Sub}^{p}_{G}$ is replaced by $\overline{{\rm Sub}^{c}_{G}}$ in the respective statement.
@@
We may note that in case $G$ is a vector group, Theorem 1.1 (1) shows that $(1)$ and $(5)$ above are not equivalent; however, in this case, statements $(2-5)$ above are equivalent by Theorem 4.1 of [40] as (2) above implies that $T\in({\rm NC})$, i.e. $\{e\}$ is not a limit point of $\{T^{n}(H)\}_{n\in{\mathbb{Z}}}$ for any closed discrete cyclic subgroup $H$ of $G$.
Recall that a semigroup $\Gamma$ of homeomorphisms of a (Hausdorff) topological space $X$ acts distally on $X$ if $\overline{\{(\gamma(x),\gamma(y))\mid\gamma\in\Gamma\}}\cap\{(d,d)\mid\ d\in X\}=\emptyset$. Note that a homeomorphism $T$ of $X$ acts distally on $X$ if and only if the cyclic group $\{T^{n}\mid n\in{\mathbb{Z}}\}$ acts distally on $X$. For a Lie group $G$, any compact subgroup of ${\rm Aut}(G)$ acts distally on ${\rm Sub}_{G}$ (see e.g. Lemma 2.2 in [39]), It follows from Theorem 1.1 of [2] that ${\mathcal{H}}$ acts distally on $G$ if and only if ${\mathcal{H}}$ acts distally on the Lie algebra of $G$, and these statements are equivalent to the statement that every element of ${\mathcal{H}}$ acts distally on $G$ (see also [11]). The following result generalises Theorems 1.1 and 1.4 to the action of any subgroup ${\mathcal{H}}$ of ${\rm Aut}(G)$ and, if ${\mathcal{H}}$ is closed, then it equates the distal action of ${\mathcal{H}}$ on ${\rm Sub}^{p}_{G}$ with that of every element of ${\mathcal{H}}$. Let $\overline{{\mathcal{H}}}$ denote the closure of ${\mathcal{H}}$ in ${\rm Aut}(G)$. It is easy to see using Abel’s results in [1, 2] that ${\mathcal{H}}$ acts distally on $G$ if and only if so does $\overline{{\mathcal{H}}}$. It follows from Ellis’ result in [16] that ${\mathcal{H}}$ acts distally on any compact ${\mathcal{H}}$-invariant subspace of ${\rm Sub}_{G}$ if and only if so does $\overline{{\mathcal{H}}}$.
@@
    Then the following hold:
    1. (a)
      $(1)$ and $(2)$ are equivalent.
    2. (b)
      $(3-6)$ are equivalent.
    3. (c)
@@
    4. (d)
      If $G={\mathbb{R}}^{d}$, for some $d\in{\mathbb{N}}$, then each of $(1)$ and $(2)$ is equivalent to the following:
@@
Note that in Theorem 1.5, one can also have equivalent statements similar to $(1-2)$, where ${\rm Sub}^{p}_{G}$ is replaced by any of $\overline{{\rm Sub}^{c}_{G}}$, ${\rm Sub}^{a}_{G}$ or ${\rm Sub}_{G}$ for the action of a subgroup ${\mathcal{H}}$. Statement (6) in the theorem implies that that ${\mathcal{H}}$ acts distally on $G$, and hence the assertion (c) above implies that Theorem 1.2 remains valid for the action of a subgroup of ${\rm Aut}(G)$ instead of an automorphism. We may note here that there is a subgroup ${\mathcal{H}}$ of ${\rm SL}(4,{\mathbb{C}})$ in which every element of ${\mathcal{H}}$ is contained in a compact subgroup of ${\rm SL}(4,{\mathbb{C}})$ but the closure of ${\mathcal{H}}$ in ${\rm SL}(4,{\mathbb{C}})$ is connected and contains unipotent elements (cf. [5]). Hence, one can not replace $\overline{{\mathcal{H}}}$ by ${\mathcal{H}}$ in (1) of Theorem 1.5.
The condition that $T$ (resp. a subgroup ${\mathcal{H}}$ of ${\rm Aut}(G)$) acts distally on the maximal compact connected central subgroup (maximal central torus) of $G$ is satisfied easily if $T\in{\rm Aut}(G)^{0}$, (resp. ${\mathcal{H}}\subset{\rm Aut}(G)^{0}$), as ${\rm Aut}(G)^{0}$ acts trivially on any compact central subgroup. It is also satisfied if $T^{n}\in{\rm Aut}(G)^{0}$ for some $n\in{\mathbb{N}}$ (resp. ${\mathcal{H}}$ or ${\mathcal{H}}{\rm Aut}(G)^{0}$ has finitely many connected components). Here, ${\rm Aut}(G)^{0}$ denotes the connected component of identity in ${\rm Aut}(G)$.
Let ${\rm Inn}(G)$ denote the group of inner automorphisms of $G$. It is a (not necessarily closed) normal subgroup of ${\rm Aut}(G)$. The following corollaries generalises the first part of Corollaries 4.5 [40].
@@
Note that ${\rm Sub}^{p}_{G}$ is contained in a larger compact set ${\rm Sub}^{co}_{G}$, which is the closure of the set of all closed connected subgroups of $G$. Hence it is easy to show that Theorems 1.11.3 are all valid if ${\rm Sub}^{p}_{G}$ is replaced by ${\rm Sub}^{co}_{G}$. In Theorems 1.4 and 1.5 and Corollary 1.6, the action of the given automorphism or a subgroup of ${\rm Aut}(G)$ is distal on ${\rm Sub}^{co}_{G}$ if and only if it is so on ${\rm Sub}^{p}_{G}$.
@@
Throughout, $G$ is a connected Lie group unless specified otherwise. Let $R$ denote the radical of $G$, i.e. the maximal connected solvable normal subgroup of $G$ and let $N$ denote the nilradical of $G$, i.e. the maximal connected nilpotent normal subgroup of $G$. Both the radical and the nilradical of $G$ are characteristic in $G$. Recall that $G$ is said to be semisimple if its radical is trivial. In particular, $G/R$ is semisimple. For a closed subgroup $H$ of $G$, let $H^{0}$ denote the connected component of the identity $e$ in $H$, $[H,H]$ denote the commutator subgroup of $H$, i.e. $[H,H]$ is generated by $\{xy,x^{-1}y^{-1}\mid x,y\in H\}$. Note that $H^{0}$, $[H,H]$ and its closure $\overline{[H,H]}$ are characteristic in $H$.
In §2, we study general properties of ${\rm Sub}_{G}$ and prove some elementary results on the structure of ${\rm Sub}^{p}_{G}$. In §3, we discuss some preliminary results on the action of ${\rm Aut}(G)$ on $G$ and ${\rm Sub}^{p}_{G}$ and prove main results stated above.
2. Actions of automorphisms of $G$ on ${\mathrm{Sub}}_{G}$ and ${\mathrm{Sub}}^{p}_{G}$
In this section, we discuss some basic properties of Chabauty topology on the space ${\rm Sub}_{G}$ of closed subgroups of $G$, the structure of ${\rm Aut}(G)$ and its action on ${\rm Sub}_{G}$. We also discuss some elementary results about the structure of ${\rm Sub}_{G}$ and ${\rm Sub}^{p}_{G}$.
Let $G$ be a locally compact Hausdorff topological group. For a compact set $K$ and an open set $U$ in $G$, we set $U_{1}(K)=\{H\in{\rm Sub}_{G}\mid H\cap K=\emptyset\}$ and $U_{2}(U)=\{H\in{\rm Sub}_{G}\mid H\cap U\neq\emptyset\}$. Here, $\{U_{1}(K)\mid K\mbox{ is compact}\}\cup\,\{U_{2}(U)\mid U\mbox{ is open}\}$ is a sub-basis of Chabauty topology on ${\rm Sub}_{G}$. As mentioned earlier, ${\rm Sub}_{G}$ is compact and Hausdorff, and if $G$ is second countable, then ${\rm Sub}_{G}$ is metrisable (cf.  [6], Lemma E.1.1). In particular, for a connected Lie group and its closed subgroups $G$, $Sub_{G}$ is metrisable.

### v2

# Characterisation of distal actions of automorphisms on the space of one-parameter subgroups of Lie groups

@@ 1. Introduction
For a Hausdorff topological space $X$, a homeomorphism $T$ of $X$ is said to be distal (equivalently, $T$ acts distally on $X$) if for any pair of distinct elements $x,y\in X$, the closure of the double orbit $\{(T^{n}(x),T^{n}(y))\mid n\in{\mathbb{Z}}\}$ in $X\times X$ does not intersect the diagonal, i.e. for $x,y\in X$ with $x\neq y$, $\overline{\{(T^{n}(x),T^{n}(y))\mid n\in{\mathbb{Z}}\}}\cap\{(a,a)\mid a\in X\}=\emptyset$. David Hilbert introduced the notion of distality to study non-ergodic actions on compact spaces (see Ellis [16] and Moore [30]). Distal actions on compact spaces, Lie groups as well as on locally compact groups have been studied by many mathematicians in different contexts (see Abels [1, 2], Choudhuri et al [10], Ellis [16], Furstenberg [18], Palit-Shah [31], Palit-Prajapati-Shah [32], Raja-Shah [34, 35], Shah [37], Shah-Yadav [39, 41, 42], and references cited therein).
@@
For a locally compact (Hausdorff) group $G$ with the identity $e$, let ${\rm Sub}_{G}$ denote the set of all closed subgroups of $G$ endowed with the Chabauty topology. Recall that ${\rm Sub}_{G}$ is compact and Hausdorff, and it is metrisable if $G$ is second countable [6]. Let ${\rm Aut}(G)$ denote the group of all automorphisms of $G$; an automorphism is both a homomorphism and a homeomorphism. There is a canonical group action of ${\rm Aut}(G)$ on ${\rm Sub}_{G}$; namely, $(T,H)\mapsto T(H)$, $T\in{\rm Aut}(G)$, $H\in{\rm Sub}_{G}$; this action gives a homomorphism from ${\rm Aut}(G)$ to ${\rm Homeo}({\rm Sub}_{G})$. As the image of ${\rm Aut}(G)$ under this action is a large subclass of homeomorphisms of ${\rm Sub}_{G}$, it is important to study the dynamics of the action of this special subclass in terms of distality.
In [42], Shah and Yadav have studied the action of automorphisms of a connected Lie groups $G$ on ${\rm Sub}_{G}$ extensively and have shown in particular that for a large class of $G$, which does not have any compact central subgroup of positive dimension, an automorphism $T$ of $G$ acts distally on ${\rm Sub}^{a}_{G}$, the space of closed abelian subgroups of $G$ if and only if $T$ generates a relatively compact subgroup in ${\rm Aut}(G)$. A similar characterisation also holds for $T$ belonging to the class (NC), i.e. those $T$ for which the closure of the $T$-orbit of any discrete cyclic group does not contain the trivial group $\{e\}$. Note that ${\rm Sub}^{a}_{G}$ is very large for many groups $G$; e.g. it is the same as ${\rm Sub}_{G}$ if $G$ is abelian. It also contains all discrete cyclic subgroups. Shah, with Palit in [31], and with Palit and Prajapati in [32], has studied distal actions of automorphisms of a discrete group $G$ on ${\rm Sub}^{a}_{G}$, where $G$ is either polycyclic or a lattice in a connected Lie group.
The question arises whether one can characterise distal actions of automorphisms on other smaller invariant subspaces of ${\rm Sub}^{a}_{G}$ for a connected Lie group $G$. Most of the techniques in earlier result mentioned above involve study of the actions on a class of discrete abelian subgroups. Here, we consider the class of smallest closed connected abelian subgroups; namely, closed one-parameter subgroups of $G$ and study the dynamics of distal actions of automorphisms on it. Let ${\rm Sub}^{p}_{G}$ denote the smallest closed subset of ${\rm Sub}_{G}$ containing all closed one-parameter subgroups of $G$. Note that ${\rm Sub}^{p}_{G}$ is compact and it is invariant under the action of ${\rm Aut}(G)$. We get a characterisation for a large class of $G$ as follows.
@@
Note that the condition on the center in Theorem 1.1 holds if $G$ is simply connected, or more generally its nilradical is simply connected. It also holds for $G/M$ for any connected Lie group $G$, where $M$ is the maximal central torus in $G$ (the maximal compact connected central subgroup in $G$). The result in the case of non-abelian groups $G$ in (2) as above gives a similar characterisation as in Theorem 4.1 of [42], even though the space ${\rm Sub}^{p}_{G}$ is much smaller than ${\rm Sub}^{a}_{G}$. In case $G$ as above is abelian, then $G={\mathbb{R}}^{d}$, a vector group, and ${\rm Sub}^{p}_{G}$ is homeomorphic to ${\mathbb{RP}}^{d-1}\sqcup\{\{0\}\}$, and the Theorem 1.1 (1) shows that a larger class of $T$ acts distally on ${\rm Sub}^{p}_{G}$.
In [42], it is shown for a connected Lie group $G$ without any compact central subgroup of positive dimension that if $T$ acts distally on ${\rm Sub}^{a}_{G}$, then it acts distally on $G$ (more generally, see Theorem 3.6 and Corollary 3.7 of [42]). We generalise this for Lie groups which are not vector groups; as in the later case, Theorem 1.1 (1) shows that it is possible to have automorphisms which act distally on ${\rm Sub}^{p}_{G}$ but not on $G$.
@@
In particular, Theorem 1.3 illustrates that the converse of Theorem 1.2 does not hold as the non-trivial unipotent automorphisms do not act distally on ${\rm Sub}^{p}_{G}$.
Let ${\rm Sub}^{c}_{G}$ denote the set of all closed discrete cyclic subgroups of $G$. It is invariant under the action of ${\rm Aut}(G)$ and so is its closure in ${\rm Sub}_{G}$. In [31] and [32], distal actions of automorphisms of ${\rm Sub}^{c}_{\Gamma}$ have been characterised, where $\Gamma$ is a certain type of discrete group, e.g. a discrete polycyclic group or a lattice in a connected Lie group. In a connected Lie group $G$, any closed one-parameter subgroup $H=\{x_{t}\}_{t\in{\mathbb{R}}}$ is the limit of $\{H_{n}\}_{n\in{\mathbb{N}}}$ as $n\to\infty$ in ${\rm Sub}_{G}$, where $H_{n}=\{x_{m/n}\mid m\in{\mathbb{Z}}\}$, $n\in{\mathbb{N}}$, are closed discrete cyclic groups. Thus ${\rm Sub}^{p}_{G}\subset\overline{{\rm Sub}^{c}_{G}}$ (the closure of ${\rm Sub}^{c}_{G}$ in ${\rm Sub}_{G})$ and the latter is contained in ${\rm Sub}^{a}_{G}$. We should note that $\overline{{\rm Sub}^{c}_{G}}$ is much bigger than ${\rm Sub}^{p}_{G}$ as the former contains all discrete cyclic subgroups (see Lemma 2.2). If $T$ acts distally on $\overline{{\rm Sub}^{c}_{G}}$ then $T\in$ (NC). It is easy to see that Theorems 1.1 (2), 1.2 and 1.3 remain valid if ${\rm Sub}^{p}_{G}$ is replaced by $\overline{{\rm Sub}^{c}_{G}}$ in the respective statements.
@@
We may note that in case $G$ is a vector group, Theorem 1.1 (1) shows that $(1)$ and $(5)$ above are not equivalent; however, in this case, statements $(2-5)$ above are equivalent by Theorem 4.1 of [42] as (2) above implies that $T\in({\rm NC})$.
Recall that a group $\Gamma$ of homeomorphisms of a (Hausdorff) topological space $X$ acts distally on $X$ if $\overline{\{(\gamma(x),\gamma(y))\mid\gamma\in\Gamma\}}\cap\{(d,d)\mid\ d\in X\}=\emptyset$ for any $x,y\in X$ with $x\neq y$. In particular, a homeomorphism $T$ of $X$ acts distally on $X$ if and only if the cyclic group $\{T^{n}\mid n\in{\mathbb{Z}}\}$ acts distally on $X$. For a Lie group $G$, any compact subgroup of ${\rm Aut}(G)$ acts distally on ${\rm Sub}_{G}$ (see e.g. Lemma 2.2 in [41]). It follows from Theorem 1.1 of [2] that for a subgroup ${\mathcal{H}}$ of ${\rm Aut}(G)$, ${\mathcal{H}}$ acts distally on $G$ if and only if ${\mathcal{H}}$ acts distally on the Lie algebra of $G$, and these statements are equivalent to the statement that every element of ${\mathcal{H}}$ acts distally on $G$ (see also [11]). Let $\overline{{\mathcal{H}}}$ denote the closure of ${\mathcal{H}}$ in ${\rm Aut}(G)$. It is easy to see using Abel’s results in [1, 2] that ${\mathcal{H}}$ acts distally on $G$ if and only if so does $\overline{{\mathcal{H}}}$. It follows from Ellis’ result in [16] that ${\mathcal{H}}$ acts distally on any compact ${\mathcal{H}}$-invariant subspace of ${\rm Sub}_{G}$ if and only if so does $\overline{{\mathcal{H}}}$.
Using Theorems 1.1 and 1.4, we get the following theorem for the action of any subgroup ${\mathcal{H}}$ of ${\rm Aut}(G)$. If ${\mathcal{H}}$ is closed, the theorem shows that ${\mathcal{H}}$ acts distally on ${\rm Sub}^{p}_{G}$ if and only if so does every element of ${\mathcal{H}}$.
@@
Then, $(1)$ and $(2)$ are equivalent and $(3-6)$ are equivalent.
@@
If $G$ is a vector group, i.e. if $G={\mathbb{R}}^{d}$, for some $d\in{\mathbb{N}}$, then $(1)$ and $(2)$ are equivalent to the following:
@@
Note that in Theorem 1.5, one can also have equivalent statements similar to $(1)$ and $(2)$, where ${\rm Sub}^{p}_{G}$ is replaced by any of $\overline{{\rm Sub}^{c}_{G}}$, ${\rm Sub}^{a}_{G}$ or ${\rm Sub}_{G}$. We may note here that one can not replace $\overline{{\mathcal{H}}}$ by ${\mathcal{H}}$ in (1) of Theorem 1.5, as there is a subgroup ${\mathcal{H}}$ of ${\rm SL}(4,{\mathbb{C}})$ in which every element of ${\mathcal{H}}$ is contained in a compact subgroup of ${\rm SL}(4,{\mathbb{C}})$ but the closure of ${\mathcal{H}}$ in ${\rm SL}(4,{\mathbb{C}})$ is connected and contains unipotent elements (cf. [5]). The condition that $T$ (resp. a subgroup ${\mathcal{H}}$ of ${\rm Aut}(G)$) acts distally on the maximal compact connected central subgroup (maximal central torus) of $G$ is satisfied easily if $T\in{\rm Aut}(G)^{0}$, (resp. ${\mathcal{H}}\subset{\rm Aut}(G)^{0}$), as ${\rm Aut}(G)^{0}$ acts trivially on any compact central subgroup. It is also satisfied if $T^{n}\in{\rm Aut}(G)^{0}$ for some $n\in{\mathbb{N}}$ (resp. ${\mathcal{H}}$ or ${\mathcal{H}}{\rm Aut}(G)^{0}$ has finitely many connected components). Here, ${\rm Aut}(G)^{0}$ denotes the connected component of the identity in ${\rm Aut}(G)$.
Let ${\rm Inn}(G)$ denote the group of inner automorphisms of $G$. It is a (not necessarily closed) normal subgroup of ${\rm Aut}(G)$. The following corollary generalises the first part of Corollary 4.5 in [42].
@@
Note that ${\rm Sub}^{p}_{G}$ is contained in a larger compact set ${\rm Sub}^{co}_{G}$, which is the closure of the set of all closed connected subgroups of $G$. Hence it is easy to show that Theorems 1.11.3 are all valid if ${\rm Sub}^{p}_{G}$ is replaced by ${\rm Sub}^{co}_{G}$. In Theorems 1.4 and 1.5 and Corollary 1.6, the action of the given automorphism or a subgroup of ${\rm Aut}(G)$ is distal on ${\rm Sub}^{co}_{G}$ if and only if it is so on ${\rm Sub}^{p}_{G}$, under the condition on the group and the action on the central torus.
@@
We now fix some notations. Let $G$ be a connected Lie group. Let $R$ (resp. $N$) denote the radical (resp. nilradical) of $G$, i.e. the largest connected solvable (resp. nilpotent) normal subgroup of $G$. Both these subgroup are characteristic in $G$. Recall that $G$ is said to be semisimple if its radical is trivial. In particular, $G/R$ is semisimple. For a closed subgroup $H$ of $G$, let $H^{0}$ denote the connected component of the identity $e$ in $H$, $[H,H]$ denote the commutator subgroup of $H$. Note that $H^{0}$, $[H,H]$ and its closure $\overline{[H,H]}$ are characteristic in $H$. Let ${\mathfrak{g}}$ denote the Lie algebra of the connected Lie group $G$. There is an exponential map $\exp:{\mathfrak{g}}\to G$ which is a continuous map and its restriction to a small open neighbourhood of 00 in ${\mathfrak{g}}$ is a homeomorphism onto an open neighbourhood of the identity $e$ in $G$. Given $T\in{\rm Aut}(G),$ there exists a corresponding Lie algebra automorphism ${\rm d\,}T:{\mathfrak{g}}\to{\mathfrak{g}}$ such that $\exp\circ\,{\rm d\,}T=T\circ\,\exp$. The map ${\rm d\,}:{\rm Aut}(G)\rightarrow{\rm Aut}({\mathfrak{g}})$, $T\mapsto{\rm d\,}T$ is an injective continuous homomorphism and its image is closed, and in fact, ${\rm d\,}$ is a homeomorphism onto its image (cf. [9], see also [22]) It is known that ${\rm Aut}({\mathfrak{g}})$ is a closed subgroup of ${\rm GL}({\mathfrak{g}})$, and we can view ${\rm Aut}(G)$ as a closed subgroup of ${\rm GL}({\mathfrak{g}})$. The topology on ${\rm Aut}(G)$, when considered as a subspace topology of ${\rm GL}({\mathfrak{g}})$, coincides with the (modified) compact-open topology (see [44]). We refer the reader to [23] for the basic structure theory of Lie groups.
In §2, we study general properties of ${\rm Sub}_{G}$ and prove some elementary results about the structure of ${\rm Sub}^{p}_{G}$. In §3, we discuss some properties and actions of automorphisms of a connected Lie group $G$ on $G$ as well as ${\rm Sub}^{p}_{G}$ and discuss some preliminary results. In §4 we prove the main results stated above.
2. Properties and structure of ${\mathrm{Sub}}_{G}$ and ${\mathrm{Sub}}^{p}_{G}$
In this section, we discuss some basic properties of Chabauty topology on the space ${\rm Sub}_{G}$. We also discuss some elementary results about the structure of ${\rm Sub}_{G}$ and ${\rm Sub}^{p}_{G}$.
Let $L$ be a locally compact Hausdorff topological group. For a compact set $K$ and an open set $U$ in $L$, let $U_{1}(K)=\{H\in{\rm Sub}_{L}\mid H\cap K=\emptyset\}$ and $U_{2}(U)=\{H\in{\rm Sub}_{L}\mid H\cap U\neq\emptyset\}$. Here, $\{U_{1}(K)\mid K\mbox{ is compact}\}\cup\,\{U_{2}(U)\mid U\mbox{ is open}\}$ is a sub-basis of Chabauty topology on ${\rm Sub}_{L}$. As mentioned earlier, ${\rm Sub}_{L}$ is compact and Hausdorff, and if $L$ is second countable, then ${\rm Sub}_{L}$ is metrisable (cf.  [6], Lemma E.1.1). In particular ${\rm Sub}_{L}$ is metrisable if $L$ is a connected Lie group. The following lemma is known (see e.g. Proposition E.1.2 in [6])